# Slamming: Training a Speech Language Model on One GPU in a Day

### v1

@@ 2 Related Work
Generative speech language models were explored under various setups (Lakhotia et al., 2021; Kharitonov et al., 2021). Lakhotia et al. (2021) were the first to show how raw and uncurated speech data can be leveraged into building a Generative Spoken Language Modeling (GSLM) system. Next, Borsos et al. (2023) proposed a cascade version using both coarse and fine speech tokens. Such a modelling framework opened up a new and promising research direction for processing and modelling spoken data, such as speech resynthesis (Polyak et al., 2021), speaking style conversion (Kreuk et al., 2021; Maimon and Adi, 2023), dialogue modelling Nguyen et al. (2022), speech-to-speech translation (Popuri et al., 2022; Peng et al., 2024b), etc. Nachmani et al. (2024) proposed augmenting a text Language Model (LM) with continuous speech data to improve spoken question-answering tasks. Recently, Park et al. (2024) proposed SLM based on state-space models (Gu et al., 2021) to further push long context-efficient modelling, while  Lin et al. (2024) proposed to fine-tune SLMs using direct preference optimisation (Rafailov et al., 2024) obtained from text LLM rankings.
@@ 3 Setup
Evaluation metrics. We assess all SLMs using four distinct evaluation metrics. The first three are based on likelihood evaluation, while the fourth is a generative metric. For likelihood based modelling we consider sBLIMP Dunbar et al. (2021), Spoken Story Cloze (sSC)), and Topic Story-Cloze (tSCHassid et al. (2024). For modelling-likelihood metrics, we evaluate the likelihood assigned by the SLMs to pairs of speech utterances, consisting of a positive example and a distractor. We calculate the percent of pairs in which the SLM assigns higher likelihood to the positive sample. sBLIMP focuses on grammatical abilities thus the negative is ungrammatical version of the positive. sSC and tSC focus on semantic modelling abilities. In sSC, the distractor suffix is taken from the original textual StoryCloze dataset (Mostafazadeh et al., 2016), allowing to assess fine-grained semantic speech understanding. In tSC, however, the distractor suffix is drawn from a different topic, enabling us to evaluate the model’s ability to understand the overall semantic concept.
Finally, to assess the generative abilities of SLMs, we compute generative perplexity (GenPPL). Following the approach of  Lakhotia et al. (2021); Hassid et al. (2024), we provide the SLM with a short speech prompt and generate speech tokens continuation. We use unit-vocoder with duration prediction to convert the tokens into speech (Polyak et al., 2021; Hassid et al., 2024). The generated speech is then transcribed, and its Perplexity (PPL) is evaluated using a pre-trained text LLM. To minimise the impact of token repetition on PPL measurements, we ground the generated text using diversity metrics derived from the auto-BLEU score (Lakhotia et al., 2021). Similarly to  Lin et al. (2024) we use bigram auto-BLEU. In other words, we ensure that all models achieve similar auto-BLEU scores, allowing for a fair comparison of PPL. Specifically, we transcribe speech segments using Whisper-large-v3333-turbo model (Radford et al., 2023) and measure PPL using Llama-$3.2$-$1$B model (LLama, 2024).
@@
To enable rapid and scalable experimentation, we developed a specialised library for SLM training that supports various model architectures, training objectives, and evaluation metrics. This library accommodates both TWIST-style training, text-speech interleaving, preference optimisation, etc. We will open-source this package along all models weights and training recipes, aiming to empower the community to further explore SLMs.
@@ 4.1 Model & Optimisation
Hyper-parameters. Unless specified otherwise, we use a context length of $512$ tokens and an effective batch size of $256$, employing gradient accumulation when necessary, as preliminary results indicated this configuration yields the best overall performance. We set the peak learning rate to $31e-$ 3 to enhance training speed and use a warmup period of $1\%$ of the total training steps, as this proved more effective than the fixed $100$-step warmup used in the original TWIST. To improve training stability, particularly with large learning rates, we apply gradient normalisation with a norm of $0.5$ at no additional cost, following Geiping and Goldstein (2023). Unless modified later in our investigation, we use an inverse-square root scheduler and the AdamW optimiser (Loshchilov, 2017).
Figure 2: Comparing PPL of different models of similar parameter count, with and without TWIST initialisation.
Initialisation. Hassid et al. (2024) empirically demonstrated that initialising SLMs with pre-trained text LMs can enhance convergence speed and improve model performance. We examine the effect of this initialisation within our setup across different model types. To do so, we train multiple models, both with and without TWIST initialisation, while staying within our compute budget. As shown in Figure 2, TWIST initialisation benefits all evaluated models at the beginning of training, though its overall impact by the end varies. Notice, the x-axis in Figure 2 represents theoretical FLOPs, calculated as $6*N_{params}*D_{tokens}$ following Hoffmann et al. (2022). However, due to variations in model architecture and implementation, practical efficiency differs, leading to varying amounts of compute processed within $24$ hours.
@@
Optimal model size & family. Cuervo and Marxer (2024) conducted a scaling analysis on GSLM-style SLMs, estimating the optimal model size and token count for a compute-efficient model. However, using a text LM initialisation might impact these findings. As we observe, TWIST initialisation greatly impact model performance, suggesting that prioritising larger models may be more effective than simply increasing the dataset size. Additionally, various model families gain different advantages from TWIST initialisation; for example, Qwen2.52.52.52.5 models show significantly better performance compared to OPT models. In Figure 3, we compare the results under the pre-defined compute budget within model families [^1] We use the text LM original names for clarity, but note that the actual size will be notably smaller due to reduced vocabulary size, e.g Qwen2.52.52.52.5-$0.5$B has $358$M parameters. Full model sizes can be found in Appendix B.. We note that the best model sizes for both MobileLLM Liu et al. (2024b), SmolLM2222 Allal et al. (2025) and Pythia Biderman et al. (2023) are $\sim 300M$ parameters, while for OPT the best is $125$M. According to Cuervo and Marxer (2024), the estimated optimal model size is approximately $66$M parameters. However, the best-performing model, Qwen2.52.52.52.5, is significantly larger. Since there are no smaller models in this family, it is difficult to determine whether this deviation is due to the quality of the initialisation or other factors. Moving forward, we proceed with both OPT-$125$M and Qwen2.52.52.52.5-$0.5$B.
@@
Positional Encoding. Transformers rely on positional encoding to capture the order of input tokens. Many modern LMs, including the Qwen models, use Rotary Position Embedding (Su et al., 2024). This method uses a hyperparameter, $\theta$, to control the trade-off between granularity and the ability to handle long contexts. $\theta$ is often tuned to accommodate longer context lengths (Yang et al., 2024a; Roziere et al., 2023). Since our context length is significantly shorter than that of the original LLM, we explore reducing $\theta$ for potential performance gains. Our findings show that setting $\theta=10,000$ with a context length of $1024$ enhances performance, so we adopt this configuration moving forward. We note that since we increase the context length, we need to reduce the batch size as well, to not run into memory problems when training. We reduce the batch size by a half and keep the same amount of gradient accumulation steps, which gives us an effective batch size of $128$.
@@
Optimiser and Scheduler. Various optimisers and schedulers have been developed to enhance training efficiency, reduce memory usage (Shazeer and Stern, 2018; Dettmers et al., 2022), or accelerate convergence (Pagliardini et al., 2024; Chen et al., 2023). With limited compute, faster convergence and better memory efficiency can be especially important. We first consider efficient optimisers, specifically AdamW with fused kernels, and $8$-bit AdamW, but observe no notable improvements in batch size or runtime compared to standard AdamW. This could do with the relatively small model size, resulting in a minimal memory footprint of the optimisers. We then compare AdamW with two state-of-the-art optimisers: AdaLomo Lv et al. (2023) and AdEMAMeix (Pagliardini et al., 2024). Results, presented in Figure 4, suggest that with the original InverseSqrt scheduler used by Hassid et al. (2024), using AdEMAMeix improves validation loss, compared to AdamW, with AdaLomo far behind.
@@ 4.2 Data
Next, we examine how the training data-mix influences performance in a compute-constrained setting. Specifically, we explore whether diversity in accents, speaking styles, etc. is beneficial and assess whether synthetic data can enhance semantic modelling abilities.
@@
Compute (GPU days) Parameters sBLIMP $\uparrow$ sStoryCloze $\uparrow$ tStoryCloze $\uparrow$ GenPPL $\downarrow$ Auto-BLEU $\downarrow$
TWIST-350M Hassid et al. (2024) 40*V100 305M 56.20 - - 137.3 3.46
@@
TWIST-7B Hassid et al. (2024) ? 7B 59.00 55.3 74.1 93.7 3.06
TWIST-13B Hassid et al. (2024) ? 13B 59.20 55.4 76.4 - -
Scaled Optimal Cuervo and Marxer (2024) ? 823M 61.3 56.7 78.0 - -
Predicted Optimal Cuervo and Marxer (2024) 1*A5000 78M 56.85 54.09 70.49 - -
@@
Table 2: Comparing slamming to leading SLMs, and predicted optimal performance for the compute. We also consider TWIST-$350$M using our code and compute budget, but with the original training recipe. $\pm$ indicates distance to min/max of $3$ seeds.
Synthetic Data. Recent studies have highlighted the potential of synthetic data generated through Text-to-Speech (TTS) (Cuervo and Marxer, 2024) or direct text-to-unit conversion (Zeng et al., 2024). Hence, we examine the impact of including synthetically generated speech within our constrained compute setup. To do so, we synthesised the TinyStories dataset (Eldan and Li, 2023) using a single-speaker TTS model (Wang et al., 2021a), as it is computationally efficient. Additionally, prior research has shown that HuBERT units largely remove speaker information (Maimon and Adi, 2023). TinyStories has been demonstrated to enhance text LM performance and improve SLMs (Cuervo and Marxer, 2024). Results are presented in Table 1. Results indicate that incorporating such synthetic data into the training data-mix significantly boosts both modelling and generative performance metrics, across all evaluated setups. We also consider adding the synthetic data to the original TWIST recipe, and the results in the bottom of Table 2 suggests that while this helps with semantic metrics, it is far from enough without other optimisations we introduced.
We see that across all datasets, and specifically with our best mixture Libri-Light, LibriSpeech and sTinyStories, Qwen-$0.5$B outperforms OPT-$125$M so we continue with it to the final stages.
@@ 4.3 Text Interleaving
For this we use Whisper-large-v3-turbo to get aligned transcriptions of our data, except sTinyStories for which we get alignment from the TTS. We follow Zeng et al. (2024) by selecting speech spans with length from a Poisson distribution with $\lambda=10$ totalling $30\%$ of the interleaved data. Following Nguyen et al. (2025) we train with balanced batches regarding token count between text data, speech data and interleaved data. We use a subset of RedPajama (Weber et al., 2024) filtered by Gopher (Rae et al., 2021) rules as our text data.
We find that under our setup performance is notably worse than speech-only Slam in all metrics. We hypothesise that in the slamming setup the larger vocabulary (leading to more parameters), and fewer speech tokens led to under-trained models. We leave for future work to find the minimal compute budget to benefit from text-interleaving or to consider other efficient approaches.
GPUs Params Num tokens sBLIMP $\uparrow$ sStoryCloze $\uparrow$ tStoryCloze $\uparrow$ GenPPL $\downarrow$ Auto-BLEU $\downarrow$
@@
GSLM Lakhotia et al. (2021) 8*V100 100M 1B 54.2 53.3 66.6 $\emptyset$ $\emptyset$
SyllableLM Baade et al. (2024) 4*A40 300M 16B 63.7 $\emptyset$ 75.4 $\emptyset$ $\emptyset$
TWIST-350M Hassid et al. (2024) 8*V100 305M 10.8B 56.20 $\emptyset$ $\emptyset$ 137.3 3.46
TWIST-1.3B Hassid et al. (2024) 32*V100 1B 10.8B 57.00 52.4 70.6 131.8 3.20
TWIST-7B Hassid et al. (2024) 32*V100 7B 36B 59.00 55.3 74.1 93.74 3.06
TWIST-13B Hassid et al. (2024) 32*V100 13B 36B 59.20 55.4 76.4 $\emptyset$ $\emptyset$
Scaled Optimal Cuervo and Marxer (2024) $\emptyset$ 823M 82B 61.3 56.7 78.0 $\emptyset$ $\emptyset$
Moshi Défossez et al. (2024) ?*H100 7B ? 58.9 58.7 81.8 $\emptyset$ $\emptyset$
SpiritLM Nguyen et al. (2025) 64*A100 7B 100B 58.0 54.8 72.9 $\emptyset$ $\emptyset$
@@
Zeng et al. (2024) $\emptyset$ 9B $\sim$1T $\emptyset$ 62.4 82.9 $\emptyset$ $\emptyset$
Moshi Défossez et al. (2024) ?*H100 7B $\sim$720B 58.8 60.8 83.0 $\emptyset$ $\emptyset$
SpiritLM Nguyen et al. (2025) 64*A100 7B 100B 58.3 61.0 82.9 $\emptyset$ $\emptyset$
AlignSLM-1.3B Lin et al. (2024) 64*A100 1B 10.8B + $\sim$158B 59.8 55.0 80.0 $\emptyset$ $\emptyset$
AlignSLM-7B Lin et al. (2024) 64*A100 7B 36B + $\sim$158B 62.3 61.1 86.8 $\emptyset$ $\emptyset$
Slam (-DPO) 2*A100 358M 16.7B 58.53 58.15 80.71 67.3 3.25
Slam 1*A5000 358M 1.4B + 5M 58.86 58.04 82.04 62.8 3.88
Slam (scaled) 2*A100 358M 16.7B + 9M 61.11 61.30 84.18 46.6 3.75
Table 3: Analysing the effect of scaling up compute for Slam. Number tokens refers to total, not necessarily unique, tokens used for training (estimated from the provided information). We separately mark DPO tokens with a +.

### v2

@@ 2 Related Work
Generative Speech Language Models were explored under various setups (Lakhotia et al., 2021; Kharitonov et al., 2021). Lakhotia et al. (2021) were the first to show how raw, uncurated speech data can be leveraged into building a Generative Speech Language Model (GSLM). Next, Borsos et al. (2023) proposed a cascade version using both coarse and fine speech tokens. Such a modelling framework opened up a new and promising research direction for processing and modelling spoken data, such as speech resynthesis (Polyak et al., 2021), speaking style conversion (Kreuk et al., 2021; Maimon and Adi, 2023), dialogue modelling Nguyen et al. (2022), speech-to-speech translation (Popuri et al., 2022; Peng et al., 2024b), etc. Nachmani et al. (2024) proposed augmenting a text Language Model (LM) with continuous speech data to improve spoken question-answering tasks. Recently, Park et al. (2024) proposed SLM based on state-space models (Gu et al., 2021) to further push long context-efficient modelling, while  Lin et al. (2024) proposed to fine-tune SLMs using direct preference optimisation (Rafailov et al., 2024) obtained from text LLM rankings.
@@ 3 Setup
Evaluation Metrics. We assess all SLMs using five distinct evaluation metrics. The first three are based on likelihood evaluation, while the fourth and fifth are generative metrics. For likelihood based modelling we consider sBLIMP Dunbar et al. (2021), Spoken Story-Cloze (sSC)), and Topic Story-Cloze (tSCHassid et al. (2024). For modelling-likelihood metrics, we evaluate the likelihood assigned by the SLMs to pairs of speech utterances, consisting of a positive example and a distractor. We calculate the percent of pairs in which the SLM assigns higher likelihood to the positive sample. sBLIMP focuses on grammatical abilities thus the negative is ungrammatical version of the positive. sSC and tSC focus on semantic modelling abilities. In sSC, the distractor suffix is taken from the original textual StoryCloze dataset (Mostafazadeh et al., 2016), allowing to assess fine-grained semantic speech understanding. In tSC, however, the distractor suffix is drawn from a different topic, enabling us to evaluate the model’s ability to understand the overall semantic concept.
To assess the generative abilities of SLMs, we compute generative perplexity (GenPPL). Following the approach of Lakhotia et al. (2021); Hassid et al. (2024), we provide the SLM with a short speech prompt and generate speech tokens continuation. We use unit-vocoder with duration prediction to convert the tokens into speech (Polyak et al., 2021; Hassid et al., 2024). The generated speech is then transcribed, and its Perplexity (PPL) is evaluated using a pre-trained text LLM. To minimise the impact of token repetition on PPL measurements, we ground the generated text using diversity metrics derived from the auto-BLEU score (Lakhotia et al., 2021). Similarly to  Lin et al. (2024) we use bigram auto-BLEU. In other words, we ensure that all models achieve similar auto-BLEU scores, allowing for a fair comparison of PPL. Specifically, we transcribe speech segments using Whisper-large-v3333-turbo model (Radford et al., 2023) and measure PPL using Llama-$3.2$-$1$B model (LLama, 2024). We calculate GenPPL on correct samples from the Spoken Story-Cloze dataset.
Finally, for our final models, we also compute GPTScore. Given a speech prompt and a generated continuation, we transcribe both and use GPT-4o to judge the quality of the continuation given the prompt, on a scale of 1 to 5. We follow the same setup and prompt as Lin et al. (2024) for the metric. We use this metric as the final form of evaluation, as it is the most costly to run.
@@
To enable rapid and scalable experimentation, we developed a specialised library for SLM training that supports various model architectures, training objectives, and evaluation metrics. It accommodates TWIST-style training, text-speech interleaving, preference optimisation, etc. We open-source this package along with all model weights and training recipes, aiming to empower the community to further explore SLMs.
@@ 4.1 Model & Optimisation
Hyper-parameters. Unless specified otherwise, we use a context length of $512$ tokens and an effective batch size of $256$, employing gradient accumulation when necessary, as preliminary results indicated this configuration yields the best overall performance. We set the peak learning rate to $31e-$ 3 to enhance training speed and use a warmup period of $1\%$ of the total training steps, as this proved more effective than the fixed $100$-step warmup used in the original TWIST. To improve training stability, particularly with large learning rates, we apply gradient normalisation with a norm of $0.5$ at no additional cost, following Geiping and Goldstein (2023). Unless modified later in our investigation, we use an inverse-square root scheduler and the AdamW optimiser (Loshchilov, 2017). Likewise, this sub-section uses the common Libri-Speech And Libri-Light datasets for training, until further investigated in Section 4.2.
Figure 2: Comparing validation PPL of different models of similar parameter count, with and without TWIST initialisation.
Initialisation. Hassid et al. (2024) empirically demonstrated that initialising SLMs with pre-trained text LMs can enhance convergence speed and improve model performance. We examine the effect of this initialisation within our setup across different model types. To do so, we train multiple models, both with and without TWIST initialisation, while staying within our compute budget. As shown in Figure 2, TWIST initialisation benefits all evaluated models at the beginning of training, though its overall impact by the end varies. Notice, the x-axis in Figure 2 represents theoretical FLOPs, calculated as $6*N_{\mathrm{params}}*D_{\mathrm{tokens}}$ following Hoffmann et al. (2022). However, due to variations in model architecture and implementation, practical efficiency differs, leading to varying amounts of compute processed within $24$ hours.
@@
Optimal Model Size & Family. Cuervo and Marxer (2024) conducted a scaling analysis on GSLM-style SLMs, estimating the optimal model size and token count for a compute-efficient model. However, using a text LM initialisation might impact these findings. As we observe, TWIST initialisation greatly impact model performance, suggesting that prioritising larger models may be more effective than simply increasing the dataset size. Additionally, various model families gain different advantages from TWIST initialisation; for example, Qwen2.52.52.52.5 models show significantly better performance compared to OPT models. In Figure 3, we compare the results under the pre-defined compute budget within model families [^1] We use the text LM original names for clarity, but note that the actual size will be notably smaller due to reduced vocabulary size, e.g Qwen2.52.52.52.5-$0.5$B has $358$M parameters. Full model sizes can be found in Appendix B.. We note that the best model sizes for MobileLLM Liu et al. (2024b), SmolLM2222 Allal et al. (2025) and Pythia Biderman et al. (2023) are $\sim 300M$ parameters, while for OPT the best is $125$M. According to Cuervo and Marxer (2024), the estimated optimal model size is approximately $66$M parameters. However, the best-performing model, Qwen2.52.52.52.5, is significantly larger. Since there are no smaller models in this family, it is difficult to determine whether this deviation is due to the quality of the initialisation or other factors. Moving forward, we proceed with both OPT-$125$M and Qwen2.52.52.52.5-$0.5$B.
@@
Positional Encoding. Transformers rely on positional encoding to capture the order of input tokens. Many modern LMs, including the Qwen models, use Rotary Position Embedding (Su et al., 2024). This method uses a hyperparameter, $\theta$, to control the trade-off between granularity and the ability to handle long contexts. $\theta$ is often tuned to accommodate longer context lengths (Yang et al., 2024a; Roziere et al., 2023). Since our context length is significantly shorter than that of the original LLM, we explore reducing $\theta$ for potential performance gains. Our findings show that setting $\theta=10,000$ with a context length of $1024$ enhances performance, so we adopt this configuration moving forward. We note that since we increase the context length (from 512 to 1024), we need to reduce the batch size as well, to not run into memory problems when training. We reduce the batch size by a half and keep the same amount of gradient accumulation steps, which gives us an effective batch size of $128$. An ablation of this adaptation is provided in Appendix D.1
@@
Optimiser and Scheduler. Various optimisers and schedulers have been developed to enhance training efficiency, reduce memory usage (Shazeer and Stern, 2018; Dettmers et al., 2022), or accelerate convergence (Pagliardini et al., 2024; Chen et al., 2023). With limited compute, these aspects become especially important. We first consider efficient optimisers, specifically AdamW with fused kernels, and $8$-bit AdamW, but observe no notable improvements in batch size or runtime compared to standard AdamW. This could do with the relatively small model size, resulting in a minimal memory footprint of the optimisers. We then compare AdamW with two state-of-the-art optimisers: AdaLomo Lv et al. (2023) and AdEMAMeix (Pagliardini et al., 2024). Results, presented in Figure 4, suggest that with the original InverseSqrt scheduler used by Hassid et al. (2024), using AdEMAMeix improves validation loss, compared to AdamW, with AdaLomo far behind.
@@ 4.2 Data
Next, we examine how the training data-mix influences performance in a compute-constrained setting. Specifically, we explore whether diversity in accents, speaking styles, etc. is beneficial and assess if synthetic data can enhance semantic abilities. We provide exact statistics for each dataset in Appendix C.
@@
[…]
TWIST-350M Hassid et al. (2024) 40*V100 305M 56.20 $\emptyset$ $\emptyset$ 137.3 3.46
@@
TWIST-7B Hassid et al. (2024) $\emptyset$ 7B 59.00 55.3 74.1 93.7 3.06
TWIST-13B Hassid et al. (2024) $\emptyset$ 13B 59.20 55.4 76.4 $\emptyset$ $\emptyset$
Scaled Optimal Cuervo and Marxer (2024) $\emptyset$ 823M 61.3 56.7 78.0 $\emptyset$ $\emptyset$
Predicted Optimal Cuervo and Marxer (2024) 1*A5000 78M 56.85 54.09 70.49 $\emptyset$ $\emptyset$
@@
Table 2: Comparing slamming to leading SLMs, and predicted optimal performance for the compute. We also consider TWIST-$350$M using our code and compute budget, but with the original training recipe. $\pm$ indicates distance to min/max of $3$ seeds. BLEU is Auto-BLEU.
Synthetic Data. Recent studies have highlighted the potential of synthetic data generated through Text-to-Speech (TTS) (Cuervo and Marxer, 2024) or direct text-to-unit conversion (Zeng et al., 2024). Hence, we examine the impact of including synthetically generated speech within our constrained compute setup. To do so, we synthesised the TinyStories dataset (Eldan and Li, 2023) using a single-speaker TTS model (Wang et al., 2021a), as it is computationally efficient. Additionally, prior research has shown that HuBERT units largely remove speaker information (Maimon and Adi, 2023). TinyStories has been demonstrated to enhance text LM performance and improve SLMs (Cuervo and Marxer, 2024). Results are presented in Table 1. Results indicate that incorporating such synthetic data into the training data-mix significantly boosts both modelling and generative performance metrics, across all evaluated setups. We also consider adding the synthetic data to the original TWIST recipe, and the results in the bottom of Table 2 suggests that while this helps with semantic metrics, it is far from enough without other optimisations we introduced. As a further ablation, we assess the performance of SLM when trained exclusively on synthetic data. Results suggest, perhaps unsurprisingly, this leads to a significant drop in performance relative to our baseline model, which uses both real and synthetic data. Specifically, the model trained only on synthetic data scores $52.35$ on sBLIMP, compared to $56.45$ for the baseline, and exhibits a notably higher validation loss on real data ($2.8$ vs. $1.65$). We observe this across all datasets, and specifically with our best mixture Libri-Light, LibriSpeech and sTinyStories, Qwen-$0.5$B outperforms OPT-$125$M so we continue with it to the final stages. These findings reinforce the importance of incorporating both real and synthetic data during training.
@@ 4.3 Text Interleaving
For this we use Whisper-large-v3333-turbo to get aligned transcriptions of our data, except sTinyStories for which we get alignment from the TTS. We follow Zeng et al. (2024) by selecting speech spans with length from a Poisson distribution with $\lambda=10$ totalling $30\%$ of the interleaved data. Following Nguyen et al. (2025) we train with balanced batches with respect to token count between text data, speech data and interleaved data. We use a subset of RedPajama (Weber et al., 2024) filtered by Gopher (Rae et al., 2021) rules as our text data.
The SLM trained with interleaving with the exact same setup as the speech only variant slightly underperformed compared to the speech only. We report results as the mean of three training runs. Specifically, it achieved tSC of 73.36 (compared to 78.01 for the speech only equivalent), sSC of 55.76 (vs 55.59) and sBLIMP of 55.71 (vs 56.45). We note that the interleaved SLM has much larger vocabulary which in turn means that the model has more parameters ($\sim 500M$ vs ${}\sim 360M$), which in turn means that each update step takes longer. For our budget the interleaved model only performed $\sim 11k$ steps vs $\sim 18k$ for speech only. Furthermore, out of all training tokens only about 40% are speech tokens in the interleaved setting. This could perhaps explain the slightly worse performance, and we leave for future work to find the minimal compute budget to benefit from text-interleaving.
[…]
@@
GSLM Lakhotia et al. (2021) 8*V100 100M 1B 54.2 53.3 66.6 $\emptyset$ $\emptyset$ $\emptyset$
SyllableLM Baade et al. (2024) 4*A40 300M 16B 63.7 $\emptyset$ 75.4 $\emptyset$ $\emptyset$ $\emptyset$
TWIST-350M Hassid et al. (2024) 8*V100 305M 10.8B 56.20 $\emptyset$ $\emptyset$ 137.3 3.46 $\emptyset$
TWIST-1.3B Hassid et al. (2024) 32*V100 1B 10.8B 57.00 52.4 70.6 131.8 3.20 1.82
TWIST-7B Hassid et al. (2024) 32*V100 7B 36B 59.00 55.3 74.1 93.74 3.06 2.71
TWIST-13B Hassid et al. (2024) 32*V100 13B 36B 59.20 55.4 76.4 $\emptyset$ $\emptyset$ $\emptyset$
Cuervo and Marxer (2024) $\emptyset$ 823M 82B 61.3 56.7 78.0 $\emptyset$ $\emptyset$ $\emptyset$
Moshi Défossez et al. (2024) ?*H100 7B $\emptyset$ 58.9 58.7 81.8 $\emptyset$ $\emptyset$ $\emptyset$
SpiritLM Nguyen et al. (2025) 64*A100 7B 100B 58.0 54.8 72.9 $\emptyset$ $\emptyset$ $\emptyset$
@@
Zeng et al. (2024) $\emptyset$ 9B $\sim$1T $\emptyset$ 62.4 82.9 $\emptyset$ $\emptyset$ $\emptyset$
Moshi Défossez et al. (2024) ?*H100 7B $\sim$720B 58.8 60.8 83.0 $\emptyset$ $\emptyset$ $\emptyset$
SpiritLM Nguyen et al. (2025) 64*A100 7B 100B 58.3 61.0 82.9 $\emptyset$ $\emptyset$ $\emptyset$
AlignSLM-1.3B Lin et al. (2024) 64*A100 1B 10.8B + $\sim$158B 59.8 55.0 80.0 $\emptyset$ $\emptyset$ 2.43
AlignSLM-7B Lin et al. (2024) 64*A100 7B 36B + $\sim$158B 62.3 61.1 86.8 $\emptyset$ $\emptyset$ 3.50
Slam (-DPO) 2*A100 358M 16.7B 58.53 58.15 80.71 67.3 3.25 $\emptyset$
Slam 1*A5000 358M 1.4B + 5M 58.86 58.04 82.04 62.8 3.88 2.09
Slam (scaled) 2*A100 358M 16.7B + 9M 61.11 61.30 84.18 46.6 3.75 2.69
Slam (large) 2*A100 1.3B 6.1B + 9M 61.43 61.52 85.30 41.2 3.89 2.79
Table 3: Analysing the effect of scaling up compute for Slam. # tokens refers to total, not unique, tokens used for training (estimated from provided information). We separately mark DPO tokens with a +. BLEU is Auto-BLEU.